OpenStudy (anonymous):

graph y-2=-2(x-5) will someone please help me with this?

OpenStudy (godlovesme):

I'll help you :) So can you first change the equation into slope-intercept form?

OpenStudy (godlovesme):

@angel04 u there??

OpenStudy (anonymous):

yes go ahead please

OpenStudy (godlovesme):

Okay =) So are you familiar with the slope-intercept form??

OpenStudy (anonymous):

y=mx+c

OpenStudy (godlovesme):

Y=mx+b ** Do u know what m and b stand for?

OpenStudy (anonymous):

m=slope and b is y intercept i think

OpenStudy (anonymous):

i am so confused right now

OpenStudy (godlovesme):

YES :D good job! So the first step will be changing the given equation into slope-intercept form

OpenStudy (anonymous):

will that be -2x+5?

OpenStudy (anonymous):

are you there?

OpenStudy (godlovesme):

U did the distribution correctly! But u forgot the y-2

OpenStudy (anonymous):

y-2-2x+5=

OpenStudy (godlovesme):

Not quite.. Y-2= -2x+5 So to isolate y, u must add 2 to both sides which will leave us with: y=-2x+7

OpenStudy (anonymous):

is the answer 0,7?

OpenStudy (godlovesme):

The answer should be a graph .. can u try first and I'll help u if u get stuck

OpenStudy (godlovesme):

(0,7) is the first point that u should plot

OpenStudy (anonymous):

y=5x?

OpenStudy (godlovesme):

Lemme help ya :D -2 is the same thing as -2/1 right?

OpenStudy (anonymous):

God loves me are you there?

OpenStudy (godlovesme):

Hmm the equation is y=-2x+7 Do u have a graph right now, I cant access the drawing tool :/

OpenStudy (godlovesme):

Yes I am

OpenStudy (godlovesme):

Come to this link and I'll show. U how to graph it http://www.twiddla.com/2079404

OpenStudy (godlovesme):

@angel04 u left??

OpenStudy (anonymous):

i'm coming to the link

OpenStudy (godlovesme):

Okay :)

OpenStudy (godlovesme):

@angel04 can you please close the question to let others know that you dobt need help? :)